40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 17 15:17:16 2016, gautam, Update, IMC, MCL Feedback MCLerror.pdfMCLtest.pngYarmCtrl.pdf
    Reply  Wed Nov 23 01:13:02 2016, gautam, Update, IMC, MCL Feedback MCL_plant.pdfOLG.pdfMC_armSpectra_X.pdfMC_armSpectra_Y.pdf
       Reply  Wed Nov 23 15:08:56 2016, rana, Update, IMC, MCL Feedback 
          Reply  Fri Feb 3 00:36:52 2017, gautam, Update, IMC, MCL Feedback - framing the problem 
             Reply  Fri Feb 3 11:40:09 2017, rana, Update, IMC, MCL Feedback - framing the problem 
                Reply  Mon Feb 6 17:03:41 2017, gautam, Update, IMC, MCL Feedback - simulink model updated mc40_v1.pdf
                   Reply  Mon Feb 6 18:20:08 2017, Koji, Update, IMC, MCL Feedback - simulink model updated 
                      Reply  Wed Feb 8 19:13:02 2017, gautam, Update, IMC, MCL Feedback - TF measurements 
                         Reply  Thu Feb 9 23:35:34 2017, gautam, Update, IMC, MCL Feedback - TF measurements mc40_v1.pdfCMboard_OLTF_fit.pdfFSSFast_OLTF_fit.pdfPCdrive_OLTF_measured.pdfdata.zip
Message ID: 12815     Entry time: Thu Feb 9 23:35:34 2017     In reply to: 12812
Author: gautam 
Type: Update 
Category: IMC 
Subject: MCL Feedback - TF measurements 

Here are the details as promised.

Attachment #1: Updated simulink model. Since I haven't actually run this model, all the TF blocks are annotated "???", but I will post an updated version once I have run the model (and fix some of the questionable aesthetic choices)

Attachment #2: Measured and fitted transfer functions from the "IN1" input (where the demodulated MC REFL goes) to the "SERVO" output of the MC servo board (to FSS box). As mentioned in my previous elog, I had to put in a pole (fitted to be at ~2MHz, called pole 9 in the plot) in order to get good agreement between fit an measurement up to 10MHz. I didn't bother fitting all the high frequency features. Both gain sliders on the MEDM screen ("IN1 Gain" and "VCO gain") were set to 0dB for this measurement, while the super boosts were all OFF.

Attachment #3: Measured and fitted transfer function from "TEST 1 IN" to "FAST OUT" of the FSS box. Both gains on the FSS MEDM screen ("Common gain adjust" and "fast gain adjust") were set to 0dB for this measurement. I didn't need any ad-hoc poles and zeros for this fit (i.e. I can map all the fitted poles and zeros to the schematic), but the fit starts to deviate from the measurement just below 1 MHz.. perhaps I need to add a zero above 1MHz, but I can't see why from the schematic...

Attachment #4: Measured TF from "TEST 1 IN" to "PC OUT" on the FSS box. MEDM gains were once again 0dB. I can't get a good fit to this, mainly because I can't decipher the poles and zeros for this path from the schematic (there are actually deviations from the schematic posted on the 40m DCC page in terms of component values, I will try and correct whatever I notice. I'll work on this...

Attachment #5: Data files + .fil files used to fit the data with LISO

Quote:

 

Data + plots + fits + updated schematics to follow...

Most of the model has come together, I am not too far from matching the modelled OLG to the measured OLG. So I will now start thinking about designing the controller for the MCL part (there are a couple of TFs that have to be measured for this path).

Attachment 1: mc40_v1.pdf  33 kB  | Hide | Hide all
mc40_v1.pdf
Attachment 2: CMboard_OLTF_fit.pdf  711 kB  | Hide | Hide all
CMboard_OLTF_fit.pdf
Attachment 3: FSSFast_OLTF_fit.pdf  765 kB  | Hide | Hide all
FSSFast_OLTF_fit.pdf
Attachment 4: PCdrive_OLTF_measured.pdf  761 kB  | Hide | Hide all
PCdrive_OLTF_measured.pdf
Attachment 5: data.zip  120 kB
ELOG V3.1.3-